OpenStudy (anonymous):

If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line? (2, 1) (1, 1) (0, 1)

OpenStudy (anonymous):

y = mx + b -1 = 2*0 + b y=2x - 1 now check (2, 1) y = 2*2 -1 = 3 doesnt work (1, 1) y = 2*1 -1 =1 works (0, 1) y = 2* 0 -1 = -1 doesnt work